# Inference-time Physics Alignment of Video Generative Models with Latent World Models

### v1

@@ 1 Introduction
State-of-the-art video generative models (magi1; sora; videopoet; lumiere; wan2025; polyak2024movie) have shown remarkable capabilities in generating visually pleasing videos. Yet, the progress in generation quality has been hindered by the limited physics understanding of these models (kang2024howfar; physicsIQ; yuan2025likephys), resulting in physically implausible video generation (bansal2024videophy; bansal2025videophy2). Ensuring physics correctness in video generation is not only crucial to increase user satisfaction, but also for reliable world modeling (lecun2022path) and downstream applications such as robotics (yang2023learning) and autonomous driving (hu2023gaia).
Prior work attributes the observed physical implausibility to the pre-training stage of video generative models (kang2024howfar), which relies on minimizing pixel or feature-level reconstruction errors. Following this hypothesis, substantial work has focused on improving pre-training or post-training of video generative models by injecting physics information (yuan2025newtongen; chefer2025videojam; li2024generativedynamics; cao2024teaching; zhang2025videorepa; zhang2025think). By contrast, another line of work assumes that physically plausible videos may be found in the manifold learned by the generative model and therefore has focused on devising inference-time methods to improve physics. This is an underexplored line of research, with two contributions relying on vision-language models (VLMs) to rewrite prompts (xue2025physt2v) and plan motion (yang2025vlipp) when a motion-controllable video generative model is available (burgert2025go). Yet, in the image generation literature, alternatives to prompt rewriting (datta2023prompt; zhang2025intricatedancepromptcomplexity) have been extensively explored to search over the generated data manifold with the goal of boosting performance at inference-time. In particular, reward models have been used to perform candidate selection with search over random seeds (ma2025inferencescaling; li2025dsearch; zhang2025classicalsearch) and to guide the model’s sampling process towards high utility generations (ye2024tfg; hemmat2023feedback; askari2024cvs; dall2025chamfer; askari2025dp; ifriqi2025erg). Despite these advances, existing work has not explored the use of off-the-shelf reward models to improve the physics plausibility of video generation at inference time.
@@
We validate the effectiveness of WMReward by achieving substantial gains in physics plausibility of video generation using two state-of-the-art video generative models: MAGI-1 (magi1), a large-scale autoregressive video generative model which generates videos chunk-by-chunk, and a large-scale holistic video latent diffusion model (vLDM). We obtain results across text-to-video (T2V), image-and-text-to-video (I2V), and video-and-text-to-video (V2V) settings. In particular, we achieve a new state of the art on the challenging PhysicsIQ benchmark (physicsIQ) with a final score of $62.0\%$, surpassing the previous best by $6.78\%$. This improvement is illustrated by the qualitative examples in Figure˜1 and further validated by an additional $11.4\%$ improvement over baselines in a human-preference study. Moreover, as shown in Figure˜1, our VJEPA-based method demonstrates a strong scaling effect with the size of the search space, and outperforms VLM-based selectors (e.g. , Qwen3-VL (yang2025qwen3)) that perform near chance level. With these results, we demonstrate the viability of using latent world models to improve the physical plausibility of video generation, paving the way toward developing reliable reward models for video generation inference-time alignment.
@@ 3 Experiments
We show the effectiveness of sampling with WMReward for improving physics plausibility across three video generation setups: image-and-text-to-video (I2V), video-and-text-to-video (V2V), and text-to-video (T2V). For all experiments, we apply WMReward to MAGI-1 24B (magi1) and to vLDM 5B111vLDM is a latent video diffusion model that utilizes a DiT backbone and jointly models spatial and temporal components at once..
@@ 3.1 Image and Multiframe-conditioned Generation
* official results from the ICCV 2025 PhysicsIQ Challenge platform are MAGI-1 I2V 37.39 (+7.62) and V2V 62.64 (+7.42).
Spatial IoU $\uparrow$ Spatio- temporal IoU $\uparrow$ Weighted Spatial IoU $\uparrow$ MSE $\downarrow$ PhysicsIQ Score $\uparrow$
@@
Table˜1 demonstrates that sampling with WMReward consistently outperforms vanilla sampling for both the vLDM and MAGI-1, and achieves state-of-the-art performance (best and second-best) across baselines and evaluation dimensions. In particular, by using the appropriate sampling strategy, $\nabla$+BoN, we achieve $62.0\%$ final PhysicsIQ score on V2V generation, which significantly outperforms the previous state-of-the-art MAGI-1 model by $6.78\%$. For I2V, sampling with WMReward surpasses the previous state-of-the-art VLIPP (yang2025vlipp) by approx. $2\%$ on PhysicsIQ.
@@ 3.3 Scaling the Search Space of WMReward
+ WMReward(BoN) $\times N\phantom{0.00}$ $\times 1.00$ 8
+ WMReward($\nabla$) $\times 5.02$N $\times 4.27$ 1
+ WMReward($\nabla$+BoN) $\times 5.02N$ $\times 4.27$ 8
@@
+ WMReward(BoN) $\times N\phantom{0.00}$ $\times 1.00$ 16
+ WMReward($\nabla$) $\times 4.96\phantom{N}$ $\times 2.07$ 01
+ WMReward($\nabla$+BoN) $\times 4.96N$ $\times 2.07$ 16
@@ 4 Related Work
Diffusion and flow-based video generative models (magi1; sora; videopoet; lumiere; wan2025) learn to generate video sequences by reversing a noising process on temporal visual data. While they achieve promising visual quality in generation, they still often produce physically implausible videos (bansal2024videophy; bansal2025videophy2; physicsIQ; yuan2025likephys). In order to mitigate this shortcoming, there has been substantial work devising pre-training and post-training approaches focused on explicitly injecting physics constraints (yuan2025newtongen), enabling motion prediction (chefer2025videojam), enhancing dynamics (li2024generativedynamics) or leveraging learned physics priors from foundation models through model finetuning or distillation (cao2024teaching; zhang2025videorepa; zhang2025think). Yet, another relatively underexplored line of work has focused on devising inference-time methods to improve physics. In particular, xue2025physt2v explore the manifold learned by the generative model through prompt rewriting and yang2025vlipp leverage VLMs to perform motion planning on top of a motion-controllable video generative model (burgert2025go).
@@
Inference-Time Alignment for Image Diffusion Models. Inference-time alignment steers a pre-trained diffusion model toward desired properties during generation without retraining. On the one hand, there are derivative-free approaches (singhal2025smc; wu2023practical; lee2025adaptive; li2024svdd; li2025dsearch). Among those, BoN selects the highest-scoring candidates at the end of the denoising process (ma2025inferencescaling). SMC-based approaches (singhal2025smc; wu2023practical) assign importance weights to multiple denoising trajectories and perform resampling during generation. Value-based importance sampling methods, such as SVDD (li2024svdd), steer a single trajectory by resampling according to the local step reward. On the other hand, we have gradient-based approaches that adjust the score function (song2020score; bansal2023universal; ye2024tfg; hemmat2023feedback; askari2024cvs; dall2025chamfer; askari2025dp), steering the denoising process toward high reward regions– resulting in e.g. , high utility or high diversity samples. These assume access to differentiable reward functions. Despite their different implementations, these methods share a similar principle of using reward signals to refine local steps or enable better global exploration.

### v2

# arXiv:2601.10553v2[cs.CV] 27 Feb 2026

@@ 1 Introduction
State-of-the-art video generative models (magi1; sora; openai2025sora2; videopoet; lumiere; wan2025; polyak2024movie) have shown remarkable capabilities in generating visually pleasing videos. Yet, the progress in generation quality has been hindered by the limited physics understanding of these models (kang2024howfar; physicsIQ; yuan2025likephys), resulting in physically implausible video generation (bansal2024videophy; bansal2025videophy2). Ensuring physics correctness in video generation is not only crucial to increase user satisfaction, but also for reliable world modeling (lecun2022path) and downstream applications such as robotics (yang2023learning) and autonomous driving (hu2023gaia).
Prior work attributes the observed physical implausibility to the pre-training stage of video generative models (kang2024howfar), which relies on minimizing pixel or feature-level reconstruction errors. Following this hypothesis, substantial work has focused on improving pre-training or post-training of video generative models by injecting physics information (yuan2025newtongen; chefer2025videojam; li2024generativedynamics; cao2024teaching; zhang2025videorepa; zhang2025think). By contrast, another line of work assumes that physically plausible videos may be found in the manifold learned by the generative model and therefore has focused on devising inference-time methods to improve physics. This is an underexplored line of research, with two contributions relying on vision-language models (VLMs) to rewrite prompts (xue2025physt2v) and plan motion (yang2025vlipp) when a motion-controllable video generative model is available (burgert2025go). However, in the image generation literature, alternatives to prompt rewriting (datta2023prompt; zhang2025intricatedancepromptcomplexity) have been extensively explored to search over the generated data manifold with the goal of boosting performance at inference-time. In particular, reward models have been used to perform candidate selection with search over random seeds (ma2025inferencescaling; li2025dsearch; zhang2025classicalsearch) and to guide the model’s sampling process towards high utility generations (ye2024tfg; hemmat2023feedback; askari2024cvs; dall2025chamfer; askari2025dp; ifriqi2025erg). Despite these advances, existing work has not explored the use of off-the-shelf reward models to improve the physics plausibility of video generation at inference time.
@@
We validate the effectiveness of WMReward by achieving substantial gains in physics plausibility of video generation using three state-of-the-art video generative models: MAGI-1 (magi1), Sora2 (openai2025sora2), and a video latent diffusion model (vLDM). These models span diverse architectures, including large-scale autoregressive generation producing videos chunk-by-chunk and holistic diffusion-based generation. We obtain results across text-to-video (T2V), image-and-text-to-video (I2V), and video-and-text-to-video (V2V) settings. In particular, we achieve a new state of the art on the challenging PhysicsIQ benchmark (physicsIQ) with a final score of $62.0\%$, surpassing the previous best by $6.78\%$. This improvement is illustrated by the qualitative examples in Figure˜1 and further validated by an additional $11.4\%$ improvement over baselines in a human-preference study. Moreover, as shown in Figure˜1, our VJEPA-based method demonstrates a strong scaling effect with the size of the search space, and outperforms VLM-based selectors (e.g. , Qwen3-VL (yang2025qwen3)) that perform near chance level. With these results, we demonstrate the viability of using latent world models to improve the physical plausibility of video generation, paving the way toward developing reliable reward models for video generation inference-time alignment.
@@ 3 Experiments
We show the effectiveness of sampling with WMReward for improving physics plausibility across three video generation setups: image-and-text-to-video (I2V), video-and-text-to-video (V2V), and text-to-video (T2V). For all experiments, we apply WMReward to MAGI-1 24B (magi1) and to vLDM 5B111vLDM is a latent video diffusion model that utilizes a DiT backbone and jointly models spatial and temporal components at once., and additionally to Sora2 (openai2025sora2) on the PhysicsIQ benchmark.
@@ 3.1 Image and Multiframe-conditioned Generation
* The official results from the ICCV 2025 PhysicsIQ Challenge platform are
MAGI-1 I2V 37.39 (+7.62) and V2V 62.64 (+7.42).
Sora2 is an API model; guidance ($\nabla$) is not available.
Spatial IoU $\uparrow$ Spatio- temp. IoU $\uparrow$ Weighted Spatial IoU $\uparrow$ MSE $\downarrow$ PhysicsIQ Score $\uparrow$
@@
Wan2.2 (wan2025) 0.358 0.120 0.217 0.008 38.26
+ VideoMAE(BoN) 0.347 0.124 0.216 0.008 37.91 (-0.35)
+ Qwen2.5-VL(BoN) 0.356 0.116 0.213 0.009 37.58 (-0.68)
+ Qwen3-VL(BoN) 0.353 0.128 0.216 0.008 38.51 (+0.25)
[…]
Sora2 (openai2025sora2) 0.363 0.187 0.231 0.007 42.30
+ VideoMAE(BoN) 0.365 0.193 0.234 0.007 42.94 (+0.64)
+ Qwen2.5-VL(BoN) 0.340 0.166 0.210 0.007 38.69 (-3.61)
+ Qwen3-VL(BoN) 0.349 0.206 0.226 0.007 42.54 (+0.24)
[…]
@@
Table˜1 demonstrates that sampling with WMReward consistently outperforms vanilla sampling for vLDM, MAGI-1, and Sora2, achieving state-of-the-art performance (best and second-best) across baselines and evaluation dimensions. In particular, by using the appropriate sampling strategy, $\nabla$+BoN, we achieve $62.0\%$ final PhysicsIQ score on V2V generation, which significantly outperforms the previous state-of-the-art MAGI-1 model by $6.78\%$. For I2V, sampling with WMReward surpasses the previous state-of-the-art Sora2 (openai2025sora2) by $4.13\%$ on PhysicsIQ.
@@ 3.3 Scaling the Search Space of WMReward
[…]
@@
[…]
@@ 4 Related Work
Diffusion and flow-based video generative models (magi1; sora; openai2025sora2; videopoet; lumiere; wan2025) learn to generate video sequences by reversing a noising process on temporal visual data. While they achieve promising visual quality in generation, they still often produce physically implausible videos (bansal2024videophy; bansal2025videophy2; physicsIQ; yuan2025likephys). In order to mitigate this shortcoming, there has been substantial work devising pre-training and post-training approaches focused on explicitly injecting physics constraints (yuan2025newtongen), enabling motion prediction (chefer2025videojam), enhancing dynamics (li2024generativedynamics) or leveraging learned physics priors from foundation models through model finetuning or distillation (cao2024teaching; zhang2025videorepa; zhang2025think). Yet, another relatively underexplored line of work has focused on devising inference-time methods to improve physics. In particular, xue2025physt2v explore the manifold learned by the generative model through prompt rewriting and yang2025vlipp leverage VLMs to perform motion planning on top of a motion-controllable video generative model (burgert2025go).
@@
Inference-Time Alignment for Image Diffusion Models. Inference-time alignment steers a pre-trained diffusion model toward desired properties during generation without retraining. On the one hand, there are derivative-free approaches (singhal2025smc; wu2023practical; lee2025adaptive; li2024svdd; li2025dsearch; jain2025diffusion; park2025steerx). Among those, BoN selects the highest-scoring candidates at the end of the denoising process (ma2025inferencescaling). SMC-based approaches (singhal2025smc; wu2023practical) assign importance weights to multiple denoising trajectories and perform resampling during generation. Value-based importance sampling methods, such as SVDD (li2024svdd), steer a single trajectory by resampling according to the local step reward. On the other hand, we have gradient-based approaches that adjust the score function (song2020score; bansal2023universal; ye2024tfg; hemmat2023feedback; askari2024cvs; dall2025chamfer; askari2025dp), steering the denoising process toward high reward regions– resulting in e.g. , high utility or high diversity samples. These assume access to differentiable reward functions. Despite their different implementations, these methods share a similar principle of using reward signals to refine local steps or enable better global exploration.